### v1

# Simulation of time crystal behavior for a few boson chiral soliton model in a ring

Abstract
We present numerical simulations for a chiral soliton model with $N=2,3$ bosons in a ring, this being a few-particle version of our previous mean-field model for a quantum time crystal OW19 . Following Syrwid, Kosior, and Sacha (SKS), the notion is that a precise position measurement of one particle can lead to spontaneous formation of a bright soliton that in a time crystal should rotate intact for at least a few revolutions around the ring SKS_EPL . In their work SKS find spontaneous formation of a soliton due to the position measurement, but quantum fluctuations cause the soliton to subsequently decay before it has a chance to perform even one revolution of the ring. Based on this they conclude that time crystal dynamics are impossible. In contrast, for our few boson chiral soliton model, and allowing for imprecise (weak) measurements of the particle position, we show that time crystal behavior is possible allowing for several revolutions of the spontaneously formed soliton around the ring. We therefore argue that our chiral soliton model can realize a quantum time crystal when weak position measurements are allowed for.
I Introduction
Since Wilczek’s seminal paper initiated the field of quantum time crystals in 2012 Wil12 , it has attracted intense interest, both fundamental and applied. His model comprised $N$ bosons with attractive interactions in a ring that was pierced by a flux tube, the key idea being that this model could yield a ground state composed of a soliton that was rotating around the ring. Although this model was subsequently shown to always yield a non-rotating ground state, the time crystal idea had taken root. Now there is active research in both continuous and discrete time crystals and for a variety of platforms including optics, ultracold atoms, metamaterials, spin systems, and both closed and open systems. For recent reviews of the area see SacZak18 ; Sacha20 ; ZalLukMon23 .
The issue undermining Wilczek’s initial proposal was nicely illustrated in a paper by Syrwid, Kosior, and Sacha (SKS) in which they simulated N-bosons in a ring that is pierced by a flux tube, and which was initiated from the N-particle ground state which shows no signatures of rotation of localization SKS_EPL . What they found was that precise position measurement of one particle could lead to the spontaneous formation of a $(N-1)$-particle soliton that is rotating, but that quantum fluctuations cause this soliton to subsequently decay before it has a chance to perform even one revolution of the ring. SKS presented results for particle numbers in the range $N=10-60$, and the lifetime $t_{c}$ of the soliton was found to decrease with $N$. We observed the same decay for a non-chiral version of our model with $N=3$, both for precise and imprecise position measurements (with an uncertainty in the position measurement), the lifetime being longer for the imprecise measurement. The conclusion of SKS remains intact that Wilczek’s model cannot show time crystal behavior SKS_EPL .
In 2019 we proposed a mean-field chiral soliton model for a quantum time crystal OW19 . This paper was followed by discussion in the literature with SKS SKS_comment -OW_lack , culminating in their paper highlighted in the previous paragraph SKS_EPL . The goal of the present paper is to examine a few boson $(N=2,3)$ limit of our previous chiral soliton model to numerically assess whether time crystal behavior is possible. In particular, we find the $N$-boson ground state, perform a position measurement of one particle, and see if spontaneous formation of the resulting $(N-1)$-boson soliton can persist over several revolutions of the ring as required for time crystal behavior. We find that for an imprecise or weak position measurement time crystal behavior is possible in our chiral soliton model, whereas for a precise measurement quantum fluctuations cause the soliton to decay in accordance with SKS SKS_EPL . We therefore argue that our chiral model can realize a quantum time crystal when weak position measurements are given due consideration.
The remainder of this paper is organized as follows: Section II describes the few boson version of our chiral soliton model along with a description of the parameters involved. The case of $N=2$ particles is covered in Sec. III and shows the ideas involved in a transparent manner. Section IV contains the main results and covers the three particle case that demonstrates that our few boson chiral soliton model can display time crystal behavior in which a two-boson soliton state can execute several revolutions of the ring. Summary and conclusions are given in Sec. V.
II Chiral soliton model
II.1 N-particle Schrödinger equation
We consider a system of $N$ scalar bosons of mass $m$ located circumferentially in a ring of radius $R$. The Schrödinger equation for the N-particle wavefunction $\psi(\theta_{1},\theta_{2},\ldots\theta_{N},t)$ is in scaled units Jackiw
$i{\partial\Psi\over\partial t}=\sum_{j=1}^{N}\left(-i{\partial\over\partial\theta_{j}}-A(\theta_{1},\theta_{2},\ldots\theta_{N})\right)^{2}\Psi+{g\over 2}\rho(\theta_{1},\theta_{2},\ldots\theta_{N})\Psi.$ (1)
Here $\theta_{j}=[-\pi,\pi],j=1,2,\ldots N$ are angular coordinates around the ring, and time is in units of $2mR^{2}/\hbar$. The scaled gauge potential is given by
$A(\theta_{1},\theta_{2},\ldots\theta_{N})=A^{(0)}+\kappa\rho(\theta_{1},\theta_{2},\ldots\theta_{N}),$ (2)
where $A^{(0)}$ represents the single-particle contribution to the gauge potential, and $\rho(\theta_{1},\theta_{2},\ldots\theta_{N})$ describes the density-dependent contribution to the gauge potential. The dimensionless parameter $\kappa$ characterizes the strength of the density-dependent chiral gauge potential, and the parameter $g$ controls the density-dependent and non-chiral contribution to the system energy. The density-dependent gauge potential is built from the combinations of the two-particle interactions between the composite bosonic particles, the two-particle interaction being denoted $\eta(x)$, with $x$ the separation between the two particles. In particular, we set
$\rho(\theta_{1},\theta_{2},\ldots\theta_{N})=\sum_{i=1}^{N}\sum_{j>i}^{N}\eta(\theta_{i}-\theta_{j}),$ (3)
where $\rho(\theta_{1},\theta_{2},\ldots\theta_{N})$ is symmetric under particle exchange by virtue of the symmetry $\eta(x)=\eta(-x)$. Since the density and corresponding density-dependent gauge potential are symmetric under particle exchange, which breaks Galilean invariance, we expect chiral dynamics to appear in this model.
II.2 Model parameters
In our previous mean-field model the chiral soliton model was derived starting from the spinor Schrödinger equation for a system of two-level bosonic atoms trapped in a ring that are dipole coupled using a laser beam carrying orbital angular momentum (OAM) characterized by the winding number $\ell$. By preparing the atoms in the appropriate dressed state a scalar field analysis may be obtained in which the atoms are subject to a vector potential of the form in Eq. (2) with $A^{(0)}={\ell\over 2}$. Moreover, the density-dependent gauge potential results from a collision-induced detuning, with the consequence that the dimensionless parameter $\kappa\propto\ell$. This means that if the winding number of the laser beam $\ell=0$, then $\kappa=0$ and there will be no chiral dynamics and no possibility for time crystal behavior. The term proportional to the dimensionless parameter $g$ describes energy-shifts due to many-body interactions.
For numerical purposes it is impractical to consider s-wave many-body interactions with $\eta(x)=\delta(x)$. Instead we here use the non-local model PriValWri23
$\eta(x)=Q\cos^{2q}(x/2),\quad q=1,2,3,\ldots,$ (4)
[…]
III Two-particle case
As a first example we consider the case of two bosons in a ring. This example is over-simplified in that if one particle is measured this leaves only one particle on the ring, so that the subsequent dynamics is linear. This example is nonetheless quite transparent and illustrates the ideas involved.
III.1 Two-particle Schrödinger equation
The density for the two bosonic particles may be written as
$\rho(\theta_{1},\theta_{2})$$\equiv$ $\eta(\theta_{1}-\theta_{2}),$ (5)
and the Schrödinger equation for $\Psi(\theta_{1},\theta_{2},t)$ then becomes PriValWri23
$i{\partial\Psi\over\partial t}=\sum_{j=1}^{2}\left(-i{\partial\over\partial\theta_{j}}-{\ell\over 2}-\kappa\eta(\theta_{1}-\theta_{2}))\right)^{2}\Psi+{g\over 2}\eta(\theta_{1}-\theta_{2})\Psi.$ (6)
Changing to center-of-mass (COM) and relative coordinates using $s=(\theta_{1}+\theta_{2})/2,x=(\theta_{1}-\theta_{2})$, the transformed Schrödinger equation for $\Psi(x,s,t)$ becomes
$i{\partial\Psi\over\partial t}=\left[-2{\partial^{2}\over\partial x^{2}}-{1\over 2}{\partial^{2}\over\partial s^{2}}+2i\left({\ell\over 2}+\kappa\eta(x)\right){\partial\over\partial s}+2\left({\ell\over 2}+\kappa\eta(x)\right)^{2}+{g\over 2}\eta(x)\right]\Psi.$ (7)
III.2 Two-particle ground state
We seek a two-particle stationary state using the ansatz PriValWri23
$\Psi(x,s,t)=e^{-i\varepsilon t+ips}\varphi(x),$ (8)
with $\varepsilon$ the scaled energy of the two-particle system, and $p$ an even integer that is the winding number associated with the center-of-mass OAM of the two particles. Substituting this ansatz into Eq. (7) yields
$\left[-{d^{2}\over dx^{2}}+\left({(p-\ell)\over 2}-\kappa\eta(x)\right)^{2}+{g\over 2}\eta(x)\right]\varphi(x)={\varepsilon\over 2}\varphi(x),$ (9)
where $\varepsilon$ is the scaled energy of the two-particle system.
Figure 1: (a) Ground state profile $|\varphi(x)|^{2}$ versus $x$ and $g$, and (b) scaled energy $\varepsilon$ versus $g$. Parameters used are $N=2,\ell=2$, $\kappa=0.2$, and $q=50$.
Figure 2: (a) Scaled energy $\varepsilon$ versus COM winding number $p$, and (b) ground state profile $|\varphi(x)|^{2}$ versus $x$. Parameters are the same as Fig. 1 along with $g=-2$.
We have solved Eq. (9) numerically for a range of parameters using the imaginary time method LehToiElo2007 applied to Eq. (7), in particular to obtain the ground state $\varphi(x)$ and associated scaled energy $\varepsilon$. Figure 1 shows numerical results for parameters $\ell=2,\kappa=0.2,q=50$. Specifically, Fig. 1(a) shows a color coded plot of the ground state profile $|\varphi(x)|^{2}$ versus $x$ for a variety of $g<0$ along the horizontal axis, and plot (b) shows the scaled energy $\varepsilon$ over the same range of $g$. The ground state profile $|\varphi(x)|^{2}$ is normalized to unity on-axis so that the profile can be discerned for each value of $g$. What is evident is that for $g>-0.5$ the ground state is uniform to a high degree, whereas for $g<-0.5$ the ground state profile is localized in terms of the relative coordinate $x$.
In the remainder of this section we choose $g=-2$ as a illustrative example, though similar results could be obtained for different combinations of parameters. Figure 2 shows some more detail of the solution for this choice. In particular, Fig. 2(a) shows the scaled energy versus $p$ and reveals that the ground state occurs for $p=\ell=2$, and (b) shows a line plot of the ground state profile $|\varphi(x)|^{2}$ (properly normalized) versus $x$ for $p=2$. Furthermore, the width (half-width at half-maximum) of the ground state profile in Fig. 2(b) with respect the the relative coordinate $x$ of the two particles is $w\sim 1$, the parameter $w$ characterizing the expected width of any localized two-particle solutions that may arise.
III.3 Time crystal dynamics from measurement
To proceed we first recognize that the ground state represented by the probability density in Fig. 2(b) does not correspond to a localized state on the ring since $|\varphi(x)|^{2}$ depends only on the relative coordinate $x$, and there is no information concerning the COM position $s$ of the particle. We can probe further and reconstruct the ground state two-body wave function $\Psi(\theta_{1},\theta_{2})$ (at $t=0$), where $|\Psi(\theta_{1},\theta_{2})|^{2}$ is shown in Fig. 3(a). From this it is straightforward to see that the probability density of finding the particle at $\theta_{1}$ for any value of $\theta_{2}$,
[…]
is a flat distribution, once again verifying that for the ground state the solution does not represent a localized solution that is pinned to a given COM position. In their work SKS SKS_EPL argued that a precise position measurement of one particle could lead initially to spontaneous formation of a localized solution, and we pursue and expand that concept here for our chiral soliton model. In particular, we assume that the position of particle $2$ is measured yielding a value $\theta_{2}^{(0)}$, but allow for the measurement to be imprecise, so that the wave function for the remaining particle may be written as
[…]
where the function $G(\theta_{2}-\theta_{2}^{(0)})$ reflects the uncertainty of the position measurement of particle $2$, and is a distribution which obeys the periodicity of the ring, and with an adjustable width $\Delta\theta$. Here for illustration we employ the following distribution
$G(\theta-\theta^{(0)})={\cal N}\cos^{2n}\left({\theta-\theta^{(0)}\over 2}\right),\quad n=1,2,3,\ldots$ (12)
with ${\cal N}$ a normalization constant. In the limit $n\gg 1$ this distribution approaches a delta-function which corresponds to the precise measurement model adopted in the paper of SKS, whereas for $n=1$ the measurement uncertainty is around $\Delta\theta\sim 2$. After the measurement, precise or imprecise, the wave function $\psi(\theta_{1},t=0)$ (suitably normalized) for the remaining particle is given by Eq. (11), and subsequent time development is governed by the single-particle Schrödinger equation
$i{\partial\psi\over\partial t}=\left(-i{\partial\over\partial\theta_{1}}-{\ell\over 2}\right)^{2}\psi.$ (13)
Figure 3(b) shows a simulation of the quantum dynamics according to Eq. (13) following a precise measurement for which we set $\theta_{2}^{(0)}=0$ and $n=50$ so that $\Delta\theta\sim 0.2$. In this case the uncertainty is much less than the expected width $w\sim 2$ for a two-particle localized state, so the measurement is expected to be very disruptive. The figure shows a color coded plot of the probability density $|\psi(\theta_{1},t)|^{2}$ versus $t$ and $\theta_{1}$. Then, for example, if we select $\theta_{1}=\theta_{2}^{(0)}=0$, and trace the probability density in time along the horizontal axis, we see what might be a periodic variation indicative of a time crystal. But this is an illusion based on us choosing $\theta_{1}=\theta_{2}^{(0)}$, a quantity we cannot know a priori, and if we chose another $\theta_{1}$ at which to monitor the time variation of the probability density we would generally find another much noisier variation with time. Another way to see this from Fig. 3(b) is that the spatial density profile $|\psi(\theta_{1},t=0)|^{2}$ becomes rapidly distorted as time increases, not simply displaced as expected for a time crystal. This is in perfect keeping with the findings of SKS that too precise a position measurement of one particle can produce a localized state initially but this will become rapidly dispersed before the initial localized solution can exhibit a single rotation around the ring.
For time crystal behavior what is required is that the center of the initial localized state basically translates in time and persists in shape for several rotations around the ring. The speed and direction of the rotation are intimately tied to the chiral gauge potential proportional to $\kappa$. Figure 4 shows an example of time crystal behavior using $n=1$ and $\Delta\theta\sim 2$, which means a much weaker measurement than in Figure 3(b). In this case the uncertainty is greater than the expected width $w\sim 1$ for a two-particle localized state, so the measurement is expected to be less disruptive. This figure shows that the initial localized solution remains largely intact while it rotates, with a velocity $v\sim-1$, and it persists over several rotations around the ring. The initial position of the centroid of the localized solution is set by the outcome of the initial measurement, $\theta^{(0)}=0$ in Fig. 4(b). In addition, if one selects a value for $\theta_{1}\neq 0$ in Fig. 4 the resulting time variation of the probability density is then largely same for any other value of $\theta_{1}$, just displaced in time. This example therefore exhibits time crystal behavior to a good approximation. This may seem counter-intuitive at first, since one would expect a localized initial state to expand and eventually interfere with itself in the ring giving rise to complicated dynamics. However, it is possible to construct special solutions to the single particle Schrödinger equation for a ring, that are localised non-dispersive solutions with a non-zero angular momentum, i.e. they rotate. Such solutions have been found by Qin Qin and are of the form
$\Phi(\theta,t)=C_{0}\cos(m\theta-nt)e^{i(m\theta-{1\over 2}(m^{2}+n^{2})t)},$ (14)
where $m$ and $n$ are integers. From Eq. (14) we notice that $\Phi(\theta,t)$ is a travelling localized solution with velocity $v=n/m$. For $m=-n=1$ we find $v=-1$, in qualitative agreement with the results in Fig. 4. In addition, because of the linearity of the single-particle Schröinger equation, we can always construct superpositions of different states such that we have a non-zero background. This is indeed the type of solution we generally find after measuring one particle with sufficiently large uncertainty. We therefore conclude that when extracting information about the two-particle ground state by measuring the position of one particle with sufficiently large uncertainty, the ground state is projected onto a single particle wave function which is localised, shape preserving, and rotating in the ring.
Figure 3: (a) Reconstructed ground state probability density $|\Psi(\theta_{1},\theta_{2})|^{2}$, and (b) simulation for a measurement with $n=50$ that localizes particle $2$ with an uncertainty $\Delta\theta\sim 0.2$. Picking a given value of $\theta_{1}$ one can trace the time evolution of the system and it does not reflect time crystal behavior. Parameters are the same as as in Fig. 1 along with $g=-2$.
Figure 4: Simulation for an imprecise measurement with $n=1$ and an uncertainty $\Delta\theta\sim 2$ that couples the remaining particle after the measurement of particle 2 to a non-dispersing solution for the ring. Picking a given value of $\theta_{1}$ one can trace the time evolution of the system and it is quite periodic, consistent with time crystal behavior over several periods. Parameters are the same as as in Fig. 1 along with $g=-2$.
The basic message from our two-particle simulations is that using a weak measurement of particle position can lead to time crystal behavior for the remaining particle on the ring. This is the case since the weak measurement can dominantly excite a non-dispersive and rotating solution on the ring, whereas the more precise position measurement excites a broader wavepacket of the ring modes that leads to the observed decay and lack of time crystal behavior. The benefit of the two-particle case is that it reveals the role played by the position measurement uncertainty in a transparent manner.
IV Three-particle case
Although the two-particle case provides some insight it does not speak to the bigger issue of whether the spontaneously formed and rotating solution after the measurement can persist even for a multi-particle state, so the rotating solution may be viewed even approximately as a rotating soliton. For this reason we next consider the three-particle case so that two particles are left on the ring after the measurement, with the possibility of exciting the two-particle analogue of a soliton.
IV.1 Three-particle Schrödinger equation
The density for the $N=3$ bosonic particles may be written as
$\rho(\theta_{1},\theta_{2},\theta_{3})\equiv\eta(\theta_{1}-\theta_{2})+\eta(\theta_{1}-\theta_{3})+\eta(\theta_{2}-\theta_{3}),$ (15)
which by construction is symmetric under exchange of any pair of particle coordinates. The Schrödinger equation for $\Psi(\theta_{1},\theta_{2},\theta_{3},t)$ may then be written as
$i{\partial\Psi\over\partial t}=\sum_{j=1}^{3}\left(-i{\partial\over\partial\theta_{j}}-{\ell\over 2}-\kappa\rho(\theta_{1},\theta_{2},\theta_{3}))\right)^{2}\Psi+{g\over 2}\rho(\theta_{1},\theta_{2},\theta_{3})\Psi.$ (16)
To proceed it is useful to use the following Jacobi coordinates Jacobi appropriate to this three-body problem
$s={1\over 3}(\theta_{1}+\theta_{2}+\theta_{3}),\quad x=(\theta_{1}-\theta_{2}),\quad y=(\theta_{1}+\theta_{2}-2\theta_{3}),$ (17)
and inverting the Jacobi coordinates we obtain
$\theta_{1}$ $=$ $s+{1\over 6}y+{1\over 2}x,$
$\theta_{2}$ $=$ $s+{1\over 6}y-{1\over 2}x,$
$\theta_{3}$ $=$ $s-{1\over 3}y.$ (18)
From these results we find
$(\theta_{1}-\theta_{2})$ $=$ $x$
$(\theta_{1}-\theta_{3})$ $=$ ${1\over 2}y+{1\over 2}x,$
$(\theta_{2}-\theta_{3})$ $=$ ${1\over 2}y-{1\over 2}x,$ (19)
so that the density $\rho$ in these Jacobi coordinates becomes
$\rho(x,y)=\eta\left({x\over 2}\right)+\eta\left({y\over 4}+{x\over 4}\right)+\eta\left({y\over 4}-{x\over 4}\right),$ (20)
which is independent of the COM coordinate $s$. The transformed Schrödinger equation for $\Psi(x,y,s,t)$ then becomes
$i{\partial\Psi\over\partial t}=-\left(2{\partial^{2}\over\partial x^{2}}+6{\partial^{2}\over\partial y^{2}}+{1\over 3}{\partial^{2}\over\partial s^{2}}\right)\Psi-2i\kappa\rho(x,y){\partial\Psi\over\partial s}+3[\kappa\rho(x,y)]^{2}\Psi+{g\over 2}\rho(x,y)\Psi.$ (21)
IV.2 Three-particle ground state
We seek a three-particle stationary state using the ansatz
$\Psi(x,y,s,t)=e^{-i\varepsilon t+ips}\varphi(x,y),$ (22)
where $p$ is the winding number associated with the center-of-mass OAM of the three particles: As a result $p$ should be an integer multiple of three. Substituting the above ansatz into Eq. (21) yields
$\varepsilon\varphi(x,y)=-\left(2{\partial^{2}\over\partial x^{2}}+6{\partial^{2}\over\partial y^{2}}\right)\varphi+3\left[{p\over 3}+\kappa\rho(x,y)\right]^{2}\varphi+{g\over 2}\rho(x,y)\varphi,$ (23)
where $\varepsilon$ is the scaled energy of the three-particle system.
We have solved Eq. (23) numerically for a range of parameters using the imaginary time method LehToiElo2007 , in particular to obtain the ground state $\varphi(x,y)$ and associated scaled energy $\varepsilon$. Figure 5 shows numerical results for parameters $\ell=3,g=-2,\kappa=0.2,q=50$. Specifically, Fig. 5(a) shows the scaled energy versus $p$ and reveals that the ground state occurs for $p=\ell=3$, and (b) shows a line plot of the ground state profile $|\varphi(x,y)|^{2}$ versus $x$ and $y$ for $p=3$.
Figure 5: (a) Scaled energy $\varepsilon$ versus COM winding number $p$, and (b) ground state profile $|\varphi(x,y)|^{2}$ for $p=\ell=3$. Parameters are $\ell=3,g=-2,\kappa=0.2,q=50$.
IV.3 Time crystal dynamics from measurement
To proceed we first recognize that the ground state represented by the probability density in Fig. 5(b) does not correspond to a localized state on the ring since $|\varphi(x,y)|^{2}$ depends only on the Jacobi coordinates $x$ and $y$, and there is no information concerning the COM position $s$ of the particle. We can also reconstruct the ground state three-body wave function for one coordinate fixed, say $\theta_{3}=0$, and $|\Psi(\theta_{1},\theta_{2},\theta_{3}=0)|^{2}$ (at $t=0$) is shown in Fig. 6(a): We note that the width of this ground state density profile is around $w\sim 1$, that provides a characteristic scale for localized three-particle states that may appear in the system.
A few words are in order about transforming from $\varphi(x,y)$ to $\Psi(\theta_{1},\theta_{2},\theta_{3})$. First we note that the Jacobi coordinates $x$ and $y$ can be re-expressed as
$x=\theta_{1}^{\prime}-\theta_{2}^{\prime},\quad y=\theta_{1}^{\prime}+\theta_{2}^{\prime},$ (24)
where $\theta_{1,2}^{\prime}=\theta_{1,2}-\theta_{3}$. This means that, for a given $\theta_{3}$, the transformation between the $\varphi(x,y)$ in the Jacobi coordinates $x$ and $y$, and $\Psi(\theta_{1}^{\prime},\theta_{2}^{\prime})$ in the coordinates $\theta_{1}^{\prime}$ and $\theta_{2}^{\prime}$ can be performed using a $45^{\circ}$ rotation. We note that at this point the wave function $\Psi(\theta_{1}^{\prime},\theta_{2}^{\prime})$ should be symmetrized if needed. One can then transform from $\Psi(\theta_{1}^{\prime},\theta_{2}^{\prime})$ to $\Psi(\theta_{1},\theta_{2},\theta_{3})$ using simple linear displacements $\theta_{1,2}=\theta_{1,2}^{\prime}+\theta_{3}$.
In their work SKS argued that a precise position measurement of one particle could lead initially to spontaneous formation of a localized solution SKS_EPL . We assume that the position of particle $3$ is measured yielding a value $\theta_{3}^{(0)}$, but allow for the measurement to be imprecise, so that the wave function for the remaining particles may be written as
[…]
where the function $G(\theta_{3}-\theta_{3}^{(0)})$ reflects the uncertainty of the position measurement, and we again use the model in Eq. (12). After the measurement the wave function $\psi(\theta_{1},\theta_{2},t=0)$ (suitably normalized) for the remaining particles is given by Eq. (25), and subsequent time development is governed by the two-particle Schrödinger equation
$i{\partial\psi\over\partial t}=-\left({\partial^{2}\over\partial\theta_{1}^{2}}+{\partial^{2}\over\partial\theta_{2}^{2}}\right)\psi+2i\left({\ell\over 2}+\kappa\rho\right)\left({\partial\over\partial\theta_{1}}+{\partial\over\partial\theta_{2}}\right)\psi+2\left({\ell\over 2}+\kappa\rho\right)^{2}\psi+{g\over 2}\rho\psi,$ (26)
where $\rho=\eta(\theta_{1}-\theta_{2})$ is the scaled density for the two remaining particles.
Figure 6(b) shows a simulation of the quantum dynamics according to Eq. (26) following a precise measurement for which we set $\theta_{3}^{(0)}=0$ and $n=50$ so that $\Delta\theta\sim 0.2$. In this case the uncertainty is much less than the expected width $w\sim 1$ for a three-particle localized state, so the measurement is expected to be very disruptive. The figure shows a color coded plot of the probability density for measuring a second particle
[…]
versus $t$ and $\theta_{1}$. Then, for example, if we select $\theta_{1}=\theta_{3}^{(0)}=0$, and track the probability density in time along the horizontal axis, we see what might be a periodic variation indicative of a time crystal. But this is an illusion based on us choosing $\theta_{1}=\theta_{3}^{(0)}$, a quantity we cannot know a priori, and if we chose another $\theta_{1}$ at which to monitor the time variation of the probability density we would generally find another variation with time. Another way to see this from Fig. 6(b) is that the spatial density profile $|\psi(\theta_{1},t=0)|^{2}$ becomes rapidly distorted as time increases. This is in perfect keeping with the findings of SKS that too precise a position measurement of one particle can produce an localized state initially but this will become rapidly dispersed before the initial localized solution can exhibit a single rotation around the ring.
For time crystal behavior what is required is that the center of the initial localized state basically translates in time and persists in shape for several rotations around the ring. The speed and direction of the rotation are intimately tied to the chiral gauge potential proportional to $\kappa$. Figure 7 shows an example of time crystal behavior using $n=1$ and $\Delta\theta\sim 2$, which means a much weaker measurement than in Figure 6(b): In this case the uncertainty is greater than the expected width $w\sim 1$ for a three-particle localized state, so the measurement is expected to be less disruptive. This figure shows that the initial localized solution remains largely intact while it rotates, and it persists over several rotations around the ring. In addition, if one selects a value for $\theta_{1}$ in Fig. 7 the resulting time variation of the probability density is then largely same for any other value of $\theta_{1}$, just displaced in time. This example therefore exhibits time crystal behavior to a good approximation.
To summarise, we deduce that the initial measurement of particle $3$ leads to spontaneous formation of a localized state that is a two-particle analogue of a mean-field chiral soliton Jackiw , and that this two-particle soliton can continue rotating whilst intact over several rotation periods. This can happen here since we considered a weak measurement as opposed to the precise measurement that leads to decay of the rotating soliton, see Fig. 6(b).
Figure 6: (a) Reconstructed ground state probability density $|\Psi(\theta_{1},\theta_{2},\theta_{3}=0)|^{2}$, and (b) simulation for a measurement with $n=50$ that localizes particle $3$ with an uncertainty $\Delta\theta\sim 0.2$. Picking a given value of $\theta_{1}$ one can trace the time evolution of the system and it does not reflect time crystal behavior. Parameters are the same as as in Fig. 5 along with $g=-2$.
Figure 7: Simulation for an imprecise measurement with $n=1$ and an uncertainty $\Delta\theta\sim 2$ that couples the remaining two particles after the measurement of particle $3$ to the two-particle limit of the chiral soliton. Picking a given value of $\theta_{1}$ one can trace the time evolution of the system and it is quite periodic, consistent with time crystal behavior over several periods. Parameters are the same as as in Fig. 5 along with $g=-2$.
V Summary and conclusions
In 2019 we proposed a mean-field chiral soliton model for a quantum time crystal OW19 , and this proposal met with some criticism in the literature SKS_EPL ,SKS_comment -OW_lack . The goal of the present paper was to examine a few boson $(N=2,3)$ limit of our previous chiral soliton model to numerically assess whether time crystal behavior is possible in the context of the type of calculation in Ref. SKS_EPL . In particular, we initialize the system in the $N$-boson ground state, perform a position measurement of one particle, and see if spontaneous formation of the resulting $(N-1)$-boson localized state (soliton for $N=2$) can persist over several revolutions of the ring as required for time crystal behavior. We find that for an imprecise or weak position measurements time crystal behavior is possible in our chiral soliton model, whereas for a precise measurement quantum fluctuations cause the soliton to decay in accordance with SKS_EPL .
Clearly the small particle numbers and parameters used here are not compatible with experiments, but are rather intended to show that time crystal behavior is not ruled out in our model. We remark that the same behavior could be seen for other combinations of parameters as long as a weak measurement is employed, with $\Delta\theta>w$. Restricting to a few bosons allowed for numerical simulations to be performed for our chiral soliton model without undue approximations, but our conclusions may still be of relevance in the more general case with large $N$. According to the Wilczek model analyzed in Ref. SKS_EPL the soliton lifetime $t_{c}$ is expected to scale with $N$. Assuming that this scaling also applies to our chiral model, if $N=3$ can show time crystal behavior then it should be expected for larger $N$ as the lifetime increases. Therefore we argue that the chiral model can realize a quantum time crystal when weak position measurements are incorporated into the scheme.
Figure 8: Simulation for an imprecise measurement with $n=1$ and an uncertainty $\Delta\theta\sim 2$ that couples the remaining particle after the measurement of particle $2$ to a linear non-dispersive wave. Picking a given value of $\theta_{1}$ one can trace the time evolution of the system and it is quite periodic, consistent with time crystal behavior over several periods. Parameters are the same as in Fig. 5 along with $g=-2$.
Finally, we discuss the issue of whether our simulations point to genuine time crystal behavior. This would require that the periodic motion seen in Fig. 7 would persist for all times for $N=3$, and we cannot be assured that this is the case based on numerical solutions. What we can say for sure is the time crystal behavior can persist over several revolutions of the ring if imprecise measurements are allowed for. Specifically, the first position measurement sets the time crystal in motion and establishes a time reference by setting the initial position of the $N=2$ soliton peak, but further measurements will be required to monitor if the time crystal continues to exhibit periodic motion. Indeed, for $N=3$ we can go one step further and ask whether time crystal behavior would persist if we performed a second imprecise measurement, of say particle $2$, acting on the two-particle state in Fig. 7. Figure 8 shows a simulation for an imprecise measurement with $n=1$ and an uncertainty $\Delta\theta\sim 2$, along with $\theta_{2}^{(0)}$ chosen consistent with the probability density for such a measurement between $[-\pi,\pi]$, that couples the remaining particle at $t=20$ to a linear non-dispersing state after the measurement of particle $2$. Then picking a given value of $\theta_{1}$ one can trace the time evolution of the system and it is quite periodic, consistent with time crystal behavior over several periods. For our $N=3$ case this validates that the time crystal behavior can survive more than one measurement. In future publications we plan to extend this initial numerical study to larger particle numbers, both numerically and analytically, and thereby move the project towards experimentally accessible parameter ranges.
VI Acknowledgements
We acknowledge helpful discussions with Erika Anderson, Stewart Lang, Joel Priestley, Gerard Valentí-Rojas, and Brian Anderson.